# Instruction Tuning for Large Language Models: A Survey

### v9

@@ Abstract
This paper surveys research works in the quickly advancing field of instruction tuning (IT), which can also be referred to as supervised fine-tuning (SFT) [^1] In this paper, unless specified otherwise, supervised fine-tuning (SFT) and instruction tuning (IT) are used interchangeably. , a crucial technique to enhance the capabilities and controllability of large language models (LLMs). Instruction tuning refers to the process of further training LLMs on a dataset consisting of (instruction, output) pairs in a supervised fashion, which bridges the gap between the next-word prediction objective of LLMs and the users’ objective of having LLMs adhere to human instructions. In this work, we make a systematic review of the literature, including the general methodology of SFT, the construction of SFT datasets, the training of SFT models, and applications to different modalities, domains and application, along with analysis on aspects that influence the outcome of SFT (e.g., generation of instruction outputs, size of the instruction dataset, etc). We also review the potential pitfalls of SFT along with criticism against it, along with efforts pointing out current deficiencies of existing strategies and suggest some avenues for fruitful research. footnotetext: Zhejiang University, Shannon.AI, Nanyang Technological University, Amazon footnotetext: Email: sy_zhang@zju.edu.cn footnotetext: Project page can be found at: https://github.com/xiaoya-li/Instruction-Tuning-Survey footnotetext: * The latest update was on Aug. 11, 2025 (Version 6).
@@ 4 Instruction Tuned LLMs
[…]
@@ 5.1 Multi-modality Datasets
[…]
@@
[…]
@@ 5.2 Multi-modality Instruction Tuning Models
[…]
@@
[…]
@@ 6 Domain-specific Instruction Tuning
[…]
@@ Appendix A Datasets
[…]

### v10

@@ Abstract
This paper surveys research works in the quickly advancing field of instruction tuning (IT), which can also be referred to as supervised fine-tuning (SFT) [^1] In this paper, unless specified otherwise, supervised fine-tuning (SFT) and instruction tuning (IT) are used interchangeably. , a crucial technique to enhance the capabilities and controllability of large language models (LLMs). Instruction tuning refers to the process of further training LLMs on a dataset consisting of (instruction, output) pairs in a supervised fashion, which bridges the gap between the next-word prediction objective of LLMs and the users’ objective of having LLMs adhere to human instructions. In this work, we make a systematic review of the literature, including the general methodology of SFT, the construction of SFT datasets, the training of SFT models, and applications to different modalities, domains and application, along with analysis on aspects that influence the outcome of SFT (e.g., generation of instruction outputs, size of the instruction dataset, etc). We also review the potential pitfalls of SFT along with criticism against it, along with efforts pointing out current deficiencies of existing strategies and suggest some avenues for fruitful research. footnotetext: Zhejiang University, Shannon.AI, Nanyang Technological University, Amazon, University of Washingtonfootnotetext: Email: sy_zhang@zju.edu.cnfootnotetext: Project page can be found at: https://github.com/xiaoya-li/Instruction-Tuning-Surveyfootnotetext: * The latest update was on Aug. 11, 2025 (Version 6).
@@ 4 Instruction Tuned LLMs
Instruction fine-tuned LLMs # Params Base Model Fine-tuning Trainset Self-build Dataset Name Size Instruct-GPT (Ouyang et al., 2022) 176B GPT-3 (Brown et al., 2020b) Yes - - BLOOMZ (Muennighoff et al., 2022)1 176B BLOOM (Scao et al., 2022) No xP3 - FLAN-T5 (Chung et al., 2022)2 11B T5 (Raffel et al., 2019) No FLAN 2021 - Alpaca (Taori et al., 2023a)3 7B LLaMA (Touvron et al., 2023a) Yes - 52K Vicuna (Chiang et al., 2023)4 13B LLaMA (Touvron et al., 2023a) Yes - 70K GPT-4-LLM (Peng et al., 2023)5 7B LLaMA (Touvron et al., 2023a) Yes - 52K Claude (Bai et al., 2022b) - - Yes - - WizardLM (Xu et al., 2023a)6 7B LLaMA (Touvron et al., 2023a) Yes Evol-Instruct 70K ChatGLM2 (Du et al., 2022)7 6B GLM (Du et al., 2022) Yes - 1.1 Tokens LIMA (Zhou et al., 2023a) 65B LLaMA (Touvron et al., 2023a) Yes - 1K OPT-IML (Iyer et al., 2022)8 175B OPT (Zhang et al., 2022a) No - - Dolly 2.0 (Conover et al., 2023a)9 12B Pythia (Biderman et al., 2023) No - 15K Falcon-Instruct (Almazrouei et al., 2023a)10 40B Falcon (Almazrouei et al., 2023b) No - - Guanaco (JosephusCheung, 2021)11 7B LLaMA (Touvron et al., 2023a) Yes - 586K Minotaur (Collective, 2023)12 15B Starcoder Plus (Li et al., 2023f) No - - Nous-Hermes (NousResearch, 2023)13 13B LLaMA (Touvron et al., 2023a) No - 300K+ TÜLU (Wang et al., 2023e)14 6.7B OPT (Zhang et al., 2022a) No Mixed - YuLan-Chat (YuLan-Chat-Team, 2023)15 13B LLaMA (Touvron et al., 2023a) Yes - 250K MOSS (Tianxiang and Xipeng, 2023)16 16B - Yes - - Airoboros (Durbin, 2023)17 13B LLaMA (Touvron et al., 2023a) Yes - - UltraLM (Ding et al., 2023a)18 13B LLaMA (Touvron et al., 2023a) Yes - -
@@ 5.1 Multi-modality Datasets
MUL-TIINSTRUCT (Xu et al., 2022)1 Image-Text 5k to 5M per task 62
PMC-VQA (Zhang et al., 2023c)2 Image-Text 227k 2
LAMM (Yin et al., 2023)3 Image-Text 186k 9
@@
Vision-Flan (Xu et al., 2024b)4 Multi-Pairs Over 1M 200+
ALLAVA (Chen et al., 2024a)5 Image-Text 1.4M 2
ShareGPT4V (Chen et al., 2023a)6 Image-Text 1.2M 2
@@ 5.2 Multi-modality Instruction Tuning Models
InstructPix2Pix (Brooks et al., 2022)1 983M I/T Stable Diffusion 983M Yes 450K
LLaVA (Liu et al., 2023b)2 13B I/T CLIP (Radford et al., 2021) 400M Yes 158K
@@
InstructBLIP (1.2B) (Dai et al., 2023)4 - I/T/V BLIP-2 (Li et al., 2023d) - No -
Otter (Li et al., 2023b)5 - I/T/V OpenFlamingo (Awadalla et al., 2023) 9B Yes 2.8M
MultiModal-GPT (Gong et al., 2023)6 - I/T/V OpenFlamingo (Awadalla et al., 2023) 9B No -
@@ 6 Domain-specific Instruction Tuning
Information extraction InstructUIE (Wang et al., 2023d)2 FlanT5 (Chung et al., 2022) 11B 1.0M
Sentiment analysis IT-MTL (Varia et al., 2022)3 T5 (Raffel et al., 2019) 220M -
Writing Writing-Alpaca-7B (Zhang et al., 2023d)4 LLaMA (Touvron et al., 2023a) 7B -
CoEdIT (Raheja et al., 2023)5 FlanT5 (Chung et al., 2022) 11B
CoPoet (Chakrabarty et al., 2022)6 T5 (Raffel et al., 2019) 11B
Medical Radiology-GPT (Liu et al., 2023c)7 Alpaca (Taori et al., 2023a) 7B 122K
ChatDoctor (Li et al., 2023j)8 LLaMA (Touvron et al., 2023a) 7B 100K
ChatGLM-Med (Wang et al., 2023a)9 ChatGLM (Du et al., 2022) 6B -
Arithmetic Goat (Liu and Low, 2023)10 LLaMA (Touvron et al., 2023a) 7B 1.0M
Code WizardCoder (Luo et al., 2023)11 StarCoder (Li et al., 2023f) 15B 78K
@@ Appendix A Datasets
Type Dataset Name # of Instances # of Lang Construction Open-source Human-Crafted UnifiedQA (Khashabi et al., 2020)1 750K En human-crafted Yes UnifiedSKG (Xie et al., 2022)3 0.8M En human-crafted Yes Natural Instructions (Honovich et al., 2022)4 193K En human-crafted Yes Super-Natural Instructions (Wang et al., 2022f)5 5M 55 Lang human-crafted Yes P3 (Sanh et al., 2021)6 12M En human-crafted Yes xP3 (Muennighoff et al., 2022)7 81M 46 Lang human-crafted Yes Flan 2021 (Longpre et al., 2023)8 4.4M En human-crafted Yes COIG (Zhang et al., 2023a)9 - - - Yes InstructGPT (Ouyang et al., 2022) 13K Multi human-crafted No Dolly (Conover et al., 2023a)22 15K En human-crafted Yes LIMA (Zhou et al., 2023a)18 1K En human-crafted Yes ChatGPT (OpenAI, 2022) - Multi human-crafted No OpenAssistant (Köpf et al., 2023)17 161,443 Multi human-crafted Yes Synthetic Data (Distillation) OIG (LAION.ai, 2023)2 43M En ChatGPT (No technique reports) Yes Unnatural Instructions (Honovich et al., 2022)10 240K En InstructGPT-Generated Yes InstructWild (Xue et al., 2023)12 104K - ChatGPT-Generated Yes Evol-Instruct / WizardLM (Xu et al., 2023a)13 52K En ChatGPT-generated Yes Alpaca (Taori et al., 2023a)14 52K En InstructGPT-generated Yes LogiCoT (Liu et al., 2023a)15 - En GPT-4-Generated Yes GPT-4-LLM (Peng et al., 2023)20 52K En&Zh GPT-4-Generated Yes Vicuna (Chiang et al., 2023) 70K En Real User-ChatGPT Conversations No Baize v1 (Conover et al., 2023b)21 111.5K En ChatGPT-Generated Yes UltraChat (Ding et al., 2023a)16 675K En&Zh GPT 3/4-Generated Yes Guanaco (JosephusCheung, 2021)19 534,530 Multi GPT (Unknown Version)-Generated Yes Orca (Mukherjee et al., 2023)23 1.5M En GPT 3.5/4-Generated Yes ShareGPT24 90K Multi Real User-ChatGPT Conversations Yes WildChat25 150K Multi Real User-ChatGPT Conversations Yes WizardCoder (Luo et al., 2023) - Code LLaMa 2-Generated No Magicoder (Wei et al., 2023b)26 75K/110K Code GPT-3.5-Generated Yes WaveCoder (Yu et al., 2023) - Code GPT 4-Generated No Phi-1 (Gunasekar et al., 2023)27 6B Tokens Code Q and A GPT-3.5-Generated Yes Phi-1.5 (Li et al., 2023i) - Code Q and A GPT-3.5-Generated No Nectar (Zhu et al., 2023a)28 183K En GPT 4-Generated Yes Synthetic Data (Self-Improvement) Self-Instruct (Wang et al., 2022c)11 52K En InstructGPT-Generated Yes Instruction Backtranslation (Li et al., 2023g) 502K En LLaMa-Generated No SPIN (Chen et al., 2024b)29 49.8K En Zephyr-Generated Yes Reasoning Data PRM800K (Wang et al., 2022c)30 800K Math human-crafted & GPT-Generated Yes O1-Journey (Li et al., 2023g)31 677 Math human-crafted & GPT-Generated Yes Self-Explore (Chen et al., 2024b) - Math GPT-Generated No MARIO (Chen et al., 2024b)32 28.8K Math human-crafted & GPT-Generated Yes MathGenie (Chen et al., 2024b) 170K Math GPT-Generated No DeepSeekMath (Chen et al., 2024b)33 120B Math human-crafted & GPT/DeepSeek-Generated Yes Compute-Optimal Sampling (Chen et al., 2024b) - Math GPT-Generated No MathScale (Chen et al., 2024b)34 2M Math GPT-Generated Yes G-LLaVA (Chen et al., 2024b)35 170K Math GPT-Generated Yes